### v2

# Binary and Ternary Quantization Can Enhance Feature Discrimination

Abstract
Quantization is widely applied in machine learning to reduce computational and storage costs for both data and models. Considering that classification tasks are fundamental to the field, it is crucial to investigate how quantization impacts classification performance. Traditional research has focused on quantization errors, assuming that larger errors generally lead to lower classification accuracy. However, this assumption lacks a solid theoretical foundation and often contradicts empirical observations. For example, despite introducing significant errors, $\{0,1\}$-binary and $\{0,\pm 1\}$-ternary quantized data have sometimes achieved classification accuracy comparable or even superior to full-precision data. To reasonably explain this phenomenon, a more accurate evaluation of classification performance is required. To achieve this, we propose a direct analysis of the feature discrimination of quantized data, instead of focusing on quantization errors. Our analysis reveals that both binary and ternary quantization can potentially enhance, rather than degrade, the feature discrimination of the original data. This finding is supported by classification experiments conducted on both synthetic and real data.
Keywords: binary quantization, ternary quantization, feature discrimination, classification
1 Introduction
Quantization has been widely applied in machine learning to simplify data storage and computation complexities, while also catering to the requirements of algorithm deployment on digital hardware. In general, this operation will lead to a decrease in classification accuracy (Baras and Dey, 1999; Hoefler et al., 2021), due to reducing the precision of data or model parameters. To achieve a balance between complexity and accuracy, it is crucial to delve into the impact of quantization on classification. Currently, the impact is mainly evaluated through quantization errors, under the premise that larger quantization errors generally lead to decreased classification accuracy (Lin et al., 2016a). However, this premise lacks a solid theoretical basis (Lin et al., 2016a), as it merely adopts the quantization principle from signal processing (Gray and Neuhoff, 1998), which primarily focuses on data reconstruction fidelity rather than classification accuracy. In practice, it appears challenging to accurately evaluate the classification performance using quantization errors.
For instance, it has been observed that some extremely low bit-width quantization methods, such as 1-bit $\{0,1\}$-binary and 2-bit $\{0,\pm 1\}$-ternary quantization, which have been successively applied in large-scale retrieval (Charikar, 2002) and deep network quantization (Qin et al., 2020; Gholami et al., 2022; Wang et al., 2023), can achieve comparable or even superior classification performance than their full-precision counterparts (Courbariaux et al., 2015; Lin et al., 2016b; Lu et al., 2023), despite suffering from high quantization errors. Apparently, the remarkable classification improvement resulting from quantization should not be attributed to the significant quantization errors. This reveals the inadequacy of quantization errors in evaluating the actual classification performance. Due to the absence of a theoretical explanation, the classification improvement induced by quantization has often been regarded as incidental and received little attention. Instead of quantization errors, in the paper we demonstrate that this intriguing phenomenon can be reasonably explained by feature discrimination. Following the Fisher’s linear discriminant analysis (Fisher, 1936), we here refer to feature discrimination as the ratio between inter-class and intra-class scatters, and evaluate the classification performance based on the rule that the higher the feature discrimination, the easier the classification. To the best of our knowledge, this is the first study that exploits feature discrimination to investigate the impact of quantization on classification, although it is more direct and reasonable than quantization errors in evaluating classification performance. The scarcity of relevant research may be attributed to the nonlinearity of the quantization operation, which substantially increases the analytical complexity of feature discrimination functions.
In the paper, it is demonstrated that the impact of the threshold-based binary and ternary quantization on feature discrimination can be analyzed, when the data are appropriately modeled using a Gaussian mixture model, with each Gaussian element representing one class of data. The Gaussian mixture model is chosen here based on two considerations. Firstly, the model has been well-established for approximating the distributions of real-world data (Torralba and Oliva, 2003; Weiss and Freeman, 2007) and their feature transformations (Wainwright and Simoncelli, 1999; Lam and Goodman, 2000). Secondly, the closure property of Gaussian distributions under linear operations can simplify the analysis of the feature discrimination function. By analyzing the discrimination across varying quantization thresholds, it is found that there exist certain quantization thresholds that can enhance the discrimination of original data, thereby yielding improved classification performance. This finding is validated through extensive classification experiments on both synthetic and real data, covering diverse modalities such as images, speech, and text.
The related works are discussed as follows. As mentioned earlier, our work should be the first to take advantage of feature discrimination to investigate the impact of quantization on classification. In the filed of signal processing, there have been a few works proposed to reduce the negative impact of quantization on signal detection or classification (Poor and Thomas, 1977; Oehler and Gray, 1995). However, these studies did not employ feature discrimination analysis, distinguishing them from our research in both methodology and outcomes. Specifically, in these studies the model design accounts for both reconstruction loss and classification loss. The classification loss is commonly modeled in various ways, such as directly minimizing the classification error on quantized data (Srinivasamurthy and Ortega, 2002), enlarging the inter-class distance between quantized data (Jana and Moulin, 2000, 2003), reducing the difference between the distributions of quantized data and original data (Baras and Dey, 1999), as well as minimizing the discrepancy in classification before and after quantization (Dogahe and Murthi, 2011). Based on these loss models, the classification performance of quantized data can only approach, but not exceed, the performance of original data (Baras and Dey, 1999).
2 Problem Formulation
In this section, we specify the feature discrimination functions for the original (non-quantized) and quantized data. Prior to this, we introduce the binary and ternary quantization functions, as well as the data modeling.
2.1 Quantization functions
The binary and ternary quantization functions are formulated as
$f_{b}(x;\tau)=\begin{cases}1,&\text{if }x>\tau\\ 0,&\text{otherwise}\end{cases}$ (1)
and
$f_{t}(x;\tau)=\begin{cases}1,&\text{if }x>\tau\\ 0,&\text{if }-\tau\leq x\leq\tau\\ -1,&\text{if }x<-\tau\end{cases}$ (2)
where the threshold parameter $\tau\in(-\infty,+\infty)$ for $f_{b}(x;\tau)$, and $\tau\in[0,+\infty)$ for $f_{t}(x;\tau)$. The two functions operate element-wise on a vector $\mathbf{x}=(x_{1},x_{2},\cdots,x_{n})^{\top}\in\mathbb{R}^{n}$, namely $f_{b}(\mathbf{x};\tau)=(f_{b}(x_{1};\tau),f_{b}(x_{2};\tau),\cdots,f_{b}(x_{n};\tau))^{\top}$ and the same applies to $f_{t}(\mathbf{x};\tau)$.
2.2 Data distributions
Throughout the work, we denote each data sample using a vector. For the sake of generality, as discussed before, we assume that the data vector randomly drawn from a class is a random vector $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{n}\}^{\top}$, with its each element $X_{i}$ following a Gaussian distribution $N(\mu_{1,i},\sigma^{2})$; and similarly, for the random vector $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}^{\top}$ drawn from another class, we suppose its each element $Y_{i}\sim N(\mu_{2,i},\sigma^{2})$, where $\mu_{2,i}\neq\mu_{1,i}$. Considering that the discrimination between the two random vectors $\mathbf{X}$ and $\mathbf{Y}$ positively correlates with the discrimination between their each pair of corresponding elements $X_{i}$ and $Y_{i}$, we propose to analyze the discrimination at the element level, specifically between $X_{i}$ and $Y_{i}$, rather than between the entire vectors, $\mathbf{X}$ and $\mathbf{Y}$. For notational convenience, without causing confusion, in the sequel we will omit the subscript "$i$" of $X_{i}$ and $Y_{i}$, and write their distributions as $X\sim N(\mu_{1},\sigma^{2})$ and $Y\sim N(\mu_{2},\sigma^{2})$, where $\mu_{1}\neq\mu_{2}$. Note that we assume here that the two variables $X$ and $Y$ share the same variance $\sigma^{2}$. This assumption is common in statistical research, as the data we intend to investigate are often drawn from the same or similar scenarios and thus exhibit similar noise levels.
To facilitate data processing, we commonly standardize variables $X$ and $Y$ using Z-score standardization, which induces specific relationships between their distributions. More precisely, in a binary classification problem, the dataset we handle is a mixture, denoted as $Z$, comprising two classes of samples drawn respectively from $X$ and $Y$. Usually, the mixture $Z$ is assumed to possess a balanced class distribution, meaning that samples are drawn from $X$ and $Y$ with equal probabilities. Under this assumption, when we perform standardization by subtracting the mean and dividing by the standard deviation for each sample in $Z$, the distributions of $X$ and $Y$ (in $Z$) will become
$\tilde{X}=\frac{X-E[Z]}{\sqrt{D[Z]}}\sim N(\tilde{\mu},\tilde{\sigma}^{2}),$ (3)
and
$\tilde{Y}=\frac{Y-E[Z]}{\sqrt{D[Z]}}\sim N(-\tilde{\mu},\tilde{\sigma}^{2}),$ (4)
with
$\tilde{\mu}=\frac{(\mu_{1}-\mu_{2})/2}{\sqrt{\sigma^{2}+\frac{1}{4}(\mu_{1}-\mu_{2})^{2}}},\quad\tilde{\sigma}^{2}=\frac{\sigma^{2}}{\sigma^{2}+\frac{1}{4}(\mu_{1}-\mu_{2})^{2}},$
where $E[Z]$ and $D[Z]$ denote the expectation and variance of $Z$, which have expressions $E[Z]=\frac{1}{2}(\mu_{1}+\mu_{2})$ and $D[Z]=\sigma^{2}+\frac{1}{4}(\mu_{1}-\mu_{2})^{2}$.
From Equations (3) and (4), it can be seen that after standardization, the two classes of variables $\tilde{X}$ and $\tilde{Y}$ still exhibit Gaussian distributions, but showcase two interesting properties: 1) their means are symmetric about zero; and 2) they have the sum of the square of the mean and the variance equal to one. In the paper, we will focus our study on such standardized data, which can be characterized as follows.
Property 1 (Distributions of two classes of standardized data)
Consider a dataset comprising two classes of data, each sampled equally from two Gaussian distributions with distinct means but identical variances. Upon standardization, the two classes of data will follow the distributions $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$, respectively, where the relationship $\mu^{2}+\sigma^{2}=1$ holds, with $\mu\in(0,1)$.
2.3 Feature discrimination
Following the Fisher’s linear discriminant rule, we define the discrimination between two classes of data as the ratio of the expected inter-class distance to the expected intra-class distance, as specified below.
Definition 2 (Discrimination between two classes of data)
For two classes of data with samples respectively drawn from the variables $X$ and $Y$, the discrimination between them is defined as
$D=\frac{E[(X_{1}-Y_{1})^{2}]}{E[(X_{1}-X_{2})^{2}]+E[(Y_{1}-Y_{2})^{2}]}\quad$ (5)
where $X_{1}$ and $X_{2}$ are i.i.d. samples of $X$, and $Y_{1}$ and $Y_{2}$ are i.i.d samples of $Y$.
In the sequel, we will utilize the above definition $D$ to denote the discrimination between original (non-quantized) data; and for the binary and ternary quantized data, as detailed below, we adopt $D_{b}$ and $D_{t}$ to represent their discrimination.
Definition 3 (Discrimination between two classes of quantized data)
Following the discrimination specified in Definition 2, the discrimination between two binary quantized data $X_{b}=f_{b}(X;\tau)$ and $Y_{b}=f_{b}(Y;\tau)$, is formulated as
$D_{b}=\frac{E[(X_{1,b}-Y_{1,b})^{2}]}{E[(X_{1,b}-X_{2,b})^{2}]+E[(Y_{1,b}-Y_{2,b})^{2}]}\quad$ (6)
where $X_{1,b}$ and $X_{2,b}$ are i.i.d. samples of $X_{b}$, and $Y_{1,b}$ and $Y_{2,b}$ are i.i.d. samples of $Y_{b}$. Similarly, the discrimination between two ternary quantized data $X_{t}=f_{t}(X;\tau)$ and $Y_{t}=f_{t}(Y;\tau)$ is expressed as
$D_{t}=\frac{E[(X_{1,t}-Y_{1,t})^{2}]}{E[(X_{1,t}-X_{2,t})^{2}]+E[(Y_{1,t}-Y_{2,t})^{2}]}\quad$ (7)
where $X_{1,t}$ and $X_{2,t}$ are i.i.d. samples of $X_{t}$, and $Y_{1,t}$ and $Y_{2,t}$ are i.i.d. samples of $Y_{t}$.
2.4 Goal
The major goal of the paper is to investigate whether there exist threshold values $\tau$ in the binary quantization $f_{b}(x;\tau)$ and the ternary quantization $f_{t}(x;\tau)$, such that the quantization can enhance the feature discrimination of original data, namely having $D_{b}>D$ and $D_{t}>D$.
3 Discrimination Analysis
3.1 Theoretical results
Theorem 4 (Binary quantization)
Consider the discrimination $D$ between two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ as specified in Property 1, as well as the discrimination $D_{b}$ between their binary quantization $X_{b}=f_{b}(X;\tau)$ and $Y_{b}=f_{b}(Y;\tau)$. We have $D_{b}>D$, if there exists a quantization threshold $\tau\in(-\infty,+\infty)$ such that
$\beta-\alpha+\frac{\mu^{2}(1-2\beta)-\mu\sqrt{\mu^{2}+4\beta(1-\beta)}}{1+\mu^{2}}>0,$ (8)
where $\alpha=\Phi\left(\frac{\tau-\mu}{\sigma}\right)$ and $\beta=\Phi\left(\frac{\tau+\mu}{\sigma}\right)$, with $\Phi(\cdot)$ denoting the cumulative distribution function of the standard normal distribution.
Theorem 5 (Ternary quantization)
Consider the discrimination $D$ between two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ as specified in Property 1, as well as the discrimination $D_{t}$ between their ternary quantization $X_{t}=f_{t}(X;\tau)$ and $Y_{t}=f_{t}(Y;\tau)$. We have $D_{t}>D$, if there exists a quantization threshold $\tau\in[0,+\infty)$ such that
$\beta-\alpha+\frac{\mu^{2}-\sqrt{\mu^{4}+8\mu^{2}\beta}}{2}>0,$ (9)
where $\alpha=\Phi\left(\frac{-\tau-\mu}{\sigma}\right)$ and $\beta=\Phi\left(\frac{-\tau+\mu}{\sigma}\right)$, with $\Phi(\cdot)$ denoting the cumulative distribution function of the standard normal distribution.
Remarks: Regarding the two theorems, there are three issues worth discussing. 1) The two theorems suggest that both binary and ternary quantization methods can indeed enhance the classification performance of original data, if there exist quantization thresholds $\tau$ that can satisfy the constraints shown in Equations (8) and (9). The following numerical analysis demonstrates that the desired threshold $\tau$ does exist, when the two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ are assigned appropriate values for $\mu$ and $\sigma$. This threshold $\tau$ can be typically determined using gradient descent methods, as detailed in Appendix B. 2) Our theoretical analysis is based on the premise that the data vectors within the same class follow Gaussian distributions in each dimension of the vectors. This condition should hold true when two classes of data are readily separable, as in this case the data points within each class should cluster tightly, allowing for Gaussian approximation. This explains the recent research findings, that the binary or ternary quantization tends to achieve comparable or superior classification performance, when handling relatively simple datasets (Courbariaux et al., 2015; Lin et al., 2016b), or distinguishable features (Lu et al., 2023). 3) The conclusion we derive in Theorem 4 for $\{0,1\}$-binary quantization also applies to another popular $\{-1,1\}$-binary quantization (Qin et al., 2020), since the Euclidean distance of the former is equivalent to the cosine distance of the latter.
3.2 Numerical validation
In this part, we conduct numerical analyses for two primary objectives. Firstly, we aim to prove the existence of the desired quantization threshold $\tau$ that holds Equations (8) and (9), namely making the left sides of the two inequalities larger than their right sides (with values equal to zero). For this purpose, we compute the values of the left sides of Equations (8) and (9), through assigning specific values to $\tau$, as well as to the two variables $X$ and $Y$’s distribution parameters $\mu$ and $\sigma^{2}$. Note that we here set $\sigma^{2}=1-\mu^{2}$, $\mu\in(0,1)$, in accordance with Property 1. In Figure 1, we examine the case that fixes $\mu=0.8$ and $\sigma^{2}=0.36$, while varying the value of $\tau$ with a step width 0.01. The results for binary quantization and ternary quantization are provided in Figures 1 (a) and (c), respectively. It can be seen that for the two quantization methods, their conditions shown in Equations (8) and (9) will hold when respectively having $\tau\in[-0.2,0.2]$ and $\tau\in[0,0.5]$. This proves the existence of the desired quantization threshold $\tau$ that can enhance feature discrimination. For limited space, we here only discuss the case of $\mu=0.8$ (and $\sigma^{2}=1-\mu^{2}$) in Figure 1. By examining different $\mu\in(0,1)$ in the same way, we can find that the quantization threshold $\tau$ that holds Equations (8) and (9), is present when $\mu\in(0.76,1)$ and $\mu\in(0.66,1)$, respectively; see Figures 8 and 9 for more evidences. This result implies two consequences. On one hand, ternary quantization has more chances to enhance feature discrimination compared to binary quantization, as the former has a broader range of $\mu$. On the other hand, the enhanced discrimination tends to be achieved when $\mu$ is sufficiently large, coupled with a correspondingly small $\sigma$, or when the discrimination between two classes of data is sufficiently high.
(a) Theoretical results for binary quantization
(b) Numerical results for binary quantization
(c) Theoretical results for ternary quantization
(d) Numerical results for ternary quantization
Figure 1: Consider two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$, with $\mu=0.8$ and $\sigma^{2}=0.36$, as specified in Property 1. The values for the left and right sides of Equations (8) and (9) are provided in (a) and (c) for binary and ternary quantization, respectively; and the discrimination $D$, $D_{b}$ and $D_{t}$ statistically estimated with Equations (5), (6) and (7) are illustrated in (b) and (d) for binary and ternary quantization, respectively.
The second goal is to verify that the quantization thresholds $\tau$ we estimate with Equations (8) and (9) in Theorems 4 and 5, can indeed enhance feature discrimination. To this end, it needs to prove that the ranges of $\tau$ derived by Equations (8) and (9), such as the ones depicted in Figures 1 (a) and (c), are consistent with the ranges we can statistically estimate by the discrimination definitions $D$, $D_{b}$ and $D_{t}$, as specified in Definitions 2 and 3. To estimate the discrimination $D$, $D_{b}$ and $D_{t}$, we randomly generate 10,000 samples from $X\sim N(0.8,0.36)$ and $Y\sim N(-0.8,0.36)$, respectively, and then statistically estimate them with Equations (5), (6) and (7), across varying $\tau$ (with a step width 0.01). The results are provided in Figures 1 (b) and (d), respectively for binary and ternary quantization. It can be seen that we have $\tau\in[-0.2,0.2]$ for $D_{b}>D$, and have $\tau\in[0,0.5]$ for $D_{t}>D$. The results coincide with the theoretical results shown in Figures 1 (a) and (c), validating Theorems 4 and 5.
4 Experiments
Through previous theoretical and numerical analyses, we have demonstrated that binary and ternary quantization can enhance feature discrimination between two classes of data, when the data vectors within each class exhibit Gaussian distributions across each dimension of their feature vectors. As enhanced feature discrimination is expected to improve classification performance, this section aims to confirm this improvement by evaluating classification performance. To see the generalizability of our theoretical findings, we evaluate classification performance on both synthetic and real data. Synthetic data can accurately adhere to the distribution conditions specified in our theoretical analysis, whereas real-world data typically cannot.
4.1 Synthetic data
4.1.1 Setting
Classification.
In this section, we mainly examine the binary classification, as it directly reflects the feature discrimination between two classes, enabling the validation of our theoretical findings. We will employ two fundamental classifiers: the $k$-nearest neighbors (KNN) algorithm (with $k=5$) (Peterson, 2009), utilizing both Euclidean and cosine distances as similarity metrics, and the support vector machine (SVM) (Cortes and Vapnik, 1995), equipped with a linear kernel.
Data generation.
We generate two classes of data by randomly and independently drawing the samples from two different random vectors $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{n}\}^{\top}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}^{\top}$, for which we set $X_{i}\sim N(\mu_{i},\sigma_{i}^{2})$ and $Y_{i}\sim N(-\mu_{i},\sigma_{i}^{2})$, with $\mu_{i}\in(-1,0)\cup(0,1)$ and $\sigma_{i}^{2}=1-\mu_{i}^{2}$, according to the data distributions specified in Property 1. Considering the fact that the features of real-world data usually exhibit sparse structures (Weiss and Freeman, 2007; Kotz et al., 2012), we further suppose that the means $\mu_{i}$ decay exponentially in magnitude, i.e. $|\mu_{i+1}|/|\mu_{i}|=e^{-\lambda}$, $\lambda\geq 0$, and set $\mu_{1}=0.8$ in the following simulation. It can be seen that with the increasing of $\lambda$, the mean’s magnitude $|\mu_{i}|$ (with $i>1$) will become smaller, indicating a smaller data element $X_{i}$ (in magnitude) and a sparser data structure. However, the data element $X_{i}$ with smaller $\mu_{i}$, is not favorable for quantization to enhance feature discrimination, as indicated by previous numerical analyses. The impact of data sparsity on quantization can be investigated by increasing the value of the parameter $\lambda$.
With the data model described above, we randomly generate two classes of data, each class containing 1000 samples. The dataset is split into two parts for training and testing, in a ratio of 4:1. Then we evaluate the KNN and SVM classification on them. The classification accuracy is determined by averaging the accuracy results obtained from repeating the data generation and classification process 100 times. The major results for KNN with Euclidean distance are presented in the main text, as shown in Figures 2 to 4; and other results, such as those for KNN with cosine distance and SVM, are given in Appendix C, specifically in Figures 10 to 13. It can be seen that the three classifiers exhibit similar performance trends. For brevity, we will focus more on the results of KNN with Euclidean distance in the subsequent discussion.
4.1.2 Results
(a) Binary data ($\lambda=0$)
(b) Binary data ($\lambda=0.01$)
(c) Binary data ($\lambda=0.1$)
(d) Binary data ($\lambda=1$)
(e) Ternary data ($\lambda=0$)
(f) Ternary data ($\lambda=0.01$)
(g) Ternary data ($\lambda=0.1$)
(h) Ternary data ($\lambda=1$)
Figure 2: KNN (Euclidean distance) classification accuracy for the 10,000-dimensional binary, ternary, and original data that are generated with the varying parameter $\lambda\in\{0,0.01,0.1,1\}$, which controls the data sparsity.
(a) 1-dim binary data
(b) 100-dim binary data
(c) 10,000-dim binary data
(d) 1-dim ternary data
(e) 100-dim ternary data
(f) 10,000-dim ternary data
Figure 3: KNN (Euclidean distance) classification accuracy for the binary, ternary, and original data generated with the parameter $\lambda=1$, and with varying dimensions $n\in\{1,100,10000\}$.
Comparison between the data with different sparsity.
In Figure 2, we investigate the classification performance for the data generated with different parameters $\lambda\in\{0,0.01,0.1,1\}$, namely with different sparsity levels. Recall that the larger the $\lambda$, the smaller the $|\mu_{i}|$, or say the smaller the data element $X_{i}$ (in magnitude). By previous analyses, the data element $X_{i}$ with smaller $|\mu_{i}|$ is not conducive to enhancing feature discrimination through quantization. Nevertheless, empirically, the negative effect does not appear to be significant. From Figure 2, it can be seen that when increasing $\lambda$ from $0.1$ to $1$, there have been quantization thresholds $\tau$ that can yield better classification performance than the original, full-precision data. In addition, it noteworthy that as $\lambda$ increases, the overall classification accuracy of original data will decrease. This decreasing trend also impacts the absolute performance of the quantized data, even though it may outperform original data.
Comparison between the data with different dimensions.
The impact of data dimensions $n\in\{1,100,10000\}$ on classification is investigated in Figure 3, where the data are generated with the exponentially decaying parameter $\lambda=1$. It can be seen that with the increasing of data dimension, the range of the quantization thresholds $\tau$ that outperform original data tends to expand, but the performance advantage declines. As previously discussed, the decline should be attributed to the data element $X_{i}$ with small means $|\mu_{i}|$, whose quantity will rise with the data dimension $n$, particularly when the decay parameter $\lambda$ of $|\mu_{i}|$ is large. To alleviate this adverse effect, it is recommended to choose a relatively smaller $\lambda$ for high-dimensional data, indicating a structure that is not overly sparse. Conversely, when the high-dimensional data is highly sparse, we should reduce its dimension to enhance the classification performance under quantization.
Comparison between binary quantization and ternary quantization.
From Figures 2 and 3, it can be seen that ternary quantization surpasses binary quantization by offering broader ranges of quantization thresholds $\tau$ that can yield higher classification accuracy than original data. This observation is consistent with our previous theoretical and numerical analyses.
Comparison between KNN and SVM.
Combining the results in Figures 2, 3, and 1013, we can say that both KNN and SVM enable quantization to improve the classification accuracy of original data, within specific ranges of quantization thresholds $\tau$. If closely examining these ranges, it can be observed that KNN often performs better when using Euclidean distance than using cosine distance. This can be attributed to the advantage of Euclidean distance over cosine distance in measuring the distance between 00 and ±$\pm 1$. Also, KNN often outperforms SVM, such as the case of $\lambda=0.1$ shown in Figures 2 and 11. This is because the support vector of SVM relies on a few data points located on the boundary between two classes, which may deteriorate during quantization. In contrast, KNN depends on the high-quality data points within each class, making it resilient to quantization noise.
(a) Classification accuracy (B)
(b) Discrimination value (B)
(c) Quantization error (B)
(d) Classification accuracy (T)
(e) Discrimination value (T)
(f) Quantization error (T)
Figure 4: KNN (Euclidean distance) classification accuracy, feature discrimination values and quantization errors are provided for the synthetic data generated with the decay parameter $\lambda=1$ and data dimension $n=1$, as well as for their binary and ternary quantization derived across different thresholds $\tau$.
Comparison between classification accuracy, feature discrimination and quantization error.
As theoretically predicted, Figure 4 illustrates that higher classification accuracy tends to be achieved at quantization thresholds $\tau$ that result in greater feature discrimination values, rather than smaller quantization errors. In particular, the variation in classification accuracy with respect to $\tau$ is more closely associated with feature discrimination values than with quantization errors. This confirms the superiority of our feature discrimination analysis in assessing the classification performance of quantized data, compared to the traditional approach of using quantization errors.
4.2 Real data
4.2.1 Setting
Datasets.
The classification is conducted on three different types of datasets, including the image datasets YaleB (Lee et al., 2005), CIFAR10 (Krizhevsky and Hinton, 2009) and ImageNet1000 (Deng et al., 2009), the speech dataset TIMIT (Fisher et al., 1986), and the text dataset Newsgroup (Lang, 1995). The datasets are briefly introduced as follows. YaleB contains face images of 38 persons, with about 64 faces per person. CIFAR10 consists of 60,000 color images from 10 different classes, with each class having 6,000 images. ImageNet1000 consists of 1000 object classes, with 1,281,167 training images, 50,000 validation images, and 100,000 test images. For the above three image datasets, we separately extract their features using Discrete Wavelet Transform (DWT), ResNet18 (He et al., 2016) and VGG16 (Simonyan and Zisserman, 2014). For ease of simulation, the resulting feature vectors are dimensionally reduced by integer multiples, leading to the sizes of 1200, 5018, and 5018 respectively. From TIMIT, as in (Mohamed et al., 2011; Hutchinson et al., 2012), we extract 39 classes of 429-dimensional phoneme features for classification, totally with 1,134,138 training samples and 58,399 test samples. Newsgroup comprises 20 categories of texts, with 11,269 samples for training, and 7,505 samples for testing. The feature dimension is reduced to 5000 by selecting the top 5000 most frequent words in the bag of words, as done in (Larochelle et al., 2012).
Classification.
To validate the broad applicability of our feature discrimination analysis results, we here examine not only binary classification but also multiclass classification. Regarding the classifiers, we employ not only the fundamental KNN and SVM which mainly involve linear operations, but also more complicated classifiers involving nonlinear operations, such as the multilayer perceptron (MLP) (Rumelhart et al., 1986) and decision trees (Quinlan, 1986). For brevity, in the main text we present the classification results of YaleB, Newsgroup, and TIMIT using KNN (with Euclidean distance) and SVM, as illustrated in Figures 5 to 7. The results for other datasets, such as CIFAR10 and ImageNet1000, and other classifiers, including KNN with cosine distance, MLP and decision trees, are provided in Appendix C, specifically in Figures 14 to 20.
For each dataset, we iterate through randomly-selected class pairs to perform binary classification. The samples for training and testing are selected according to the default settings of the datasets. For YaleB without prior settings, we randomly assign half of the samples for training and the remaining half for testing. In the simulation, we need to test the classification performance of quantized data across varying quantization threshold $\tau$ values. For ease of analysis, we apply a uniform threshold $\tau$ value to all dimensions of feature vectors in each trial, though individual thresholds per dimension could potentially yield higher feature discrimination. To tackle the scale varying of the threshold $\tau$ values across different data, we here suppose $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements of all feature vectors participating in classification, and $\gamma$ is a scaling parameter. By adjusting the value of $\gamma$ within a narrow range, as illustrated later, we can derive the desired threshold $\tau$ values for various types of data.
4.2.2 Results
(a) Binary data (KNN)
(b) Binary data (SVM)
(c) Ternary data (KNN)
(d) Ternary data (SVM)
Figure 5: Classification accuracy for the binary, ternary, and original data by KNN (Euclidean distance) and SVM on YaleB. The parameter $\gamma$ corresponds to a threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors.
(a) Binary data (KNN)
(b) Binary data (SVM)
(c) Ternary data (KNN)
(d) Ternary data (SVM)
Figure 6: Classification accuracy for the binary, ternary, and original data by KNN (Euclidean distance) and SVM on TIMIT. The parameter $\gamma$ corresponds to a threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors.
(a) Binary data (KNN)
(b) Binary data (SVM)
(c) Ternary data (KNN)
(d) Ternary data (SVM)
Figure 7: Classification accuracy for the binary, ternary, and original data by KNN (Euclidean distance) and SVM on Newsgroup. The parameter $\gamma$ corresponds to a threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors.
Binary classification.
In Figures 5-7,14, 15 and 19, we present the binary classification results of KNN and SVM on five distinct datasets. It can be seen that there indeed exist specific ranges of quantization threshold $\tau$ values that allow both binary and ternary quantization to achieve superior, or comparable classification performance compared to the original data. Note that as illustrated in Figure 16, each data category approximately follows, rather than perfectly conforms to the Gaussian distribution assumption made in our theoretical analysis. This underscores the robustness of our theoretical results. Similar to the classification of synthetic data, the classification of real data exhibits the following performance trends. 1) When using Euclidean distance, KNN consistently obtains quantization threshold $\tau$ values that raise the classification accuracy of original data across all datasets. 2) In contrast to cosine distance, Euclidean distance enables KNN to cover a wider range of $\tau$ values that facilitate classification improvement. 3) With SVM, quantization occasionally achieves comparable performance to the original data, rather than surpassing it, as exemplified in Figure 14. 4) Ternary quantization often outperforms binary quantization by offering a broader range of threshold $\tau$ values conductive to classification improvement. The reasoning behind these results has been elaborated in our previous classification analysis of synthetic data. Overall, the consistency in classification performance between real data and synthetic data substantiates the rationality and generalizability of our theoretical findings.
KNN and SVM are well-suited for our linear discrimination analysis since they use linear operations to assess feature similarity. Other, more complex nonlinear classifiers, like MLP and decision trees, can also obtain the threshold $\tau$ values that enhance classification performance, as demonstrated in Figures 17 and 18. This capability stems from the fact that these nonlinear classifiers are grounded in linear operations, which serve to measure the linear similarity between features or model parameters, thereby aligning with our analytical framework.
Multiclass classification.
By enhancing feature discrimination between two classes, the performance of multiclass classification is expected to improve as well. This improvement has been observed in recent experiments (Lu et al., 2023), particularly in the challenging 1,000-class classification on deep features of ImageNet1000, as illustrated in Figure 20. In multiclass classification, we simply apply a uniform threshold to each dimension of the feature vectors. This approach conforms to our analytical framework rooted in binary classification, as classification in each dimension can be approximately considered a binary classification problem, where the values of feature elements across different classes indicate the intensity of the feature attribute at that dimension, i.e., whether it is strong or weak. The successful adaption to multiclass classification highlights the broad applicability of our theoretical findings.
5 Conclusion
The paper proposes a novel methodology that leverages feature discrimination to evaluate the influence of quantization on classification accuracy. Unlike the conventional approach, which focuses primarily on quantization errors, this feature discrimination-based analysis provides a more direct and logically sound assessment of classification performance.
The theoretical analysis reveals that common binary and ternary quantization techniques can enhance the feature discrimination of the original data, particularly when data vectors within the same class exhibit Gaussian distributions along each dimension. These theoretical insights are supported by numerical simulations. Moreover, the enhanced feature discrimination is validated through classification tasks conducted on both synthetic data adhering to Gaussian distributions and a variety of real-world datasets that do not strictly conform to Gaussian distributions, such as convolutional features extracted from images, spectral features derived from speech, and term frequency features obtained from texts. This underscores the robustness and generalizability of the theoretical findings.
A significant contribution of this research lies in its challenge to the widely accepted notion that larger quantization errors generally result in poorer classification performance. By challenging this conventional view, the study opens avenues for the exploration of more advanced quantization techniques that could potentially enhance, rather than impair, model performance. Consequently, this research is expected to drive advancements in the fields of model quantization and compression, particularly within the realms of deep neural networks (Qin et al., 2020; Gholami et al., 2022) and biological neural networks (Haufe et al., 2010; Dasgupta et al., 2017).
Appendix A Proofs for Section 3
A.1 Proof of Theorem 4
Let $X_{1}$ and $X_{2}$ be i.i.d. samples of $X$, and $Y_{1}$ and $Y_{2}$ be i.i.d. samples of $Y$. Denote $X_{i,b}$ and $Y_{i,b}$ as the binary quantization of $X_{i}$ and $Y_{i}$, i.e. $X_{i,b}=f_{b}(X_{i};\tau)$ and $Y_{i,b}=f_{b}(Y_{i};\tau)$, where $i=1,2$. By the distributions of $X$ and $Y$ specified in Property 1 and the binary quantization function $f_{b}(x;\tau)$ defined in Equation(1), the probability mass functions of $X_{i,b}$ and $Y_{i,b}$ can be derived as
$P(X_{i,b}=k)=\left\{\begin{aligned} 1&-\alpha,&\quad k=1\\ &\alpha,&\quad k=0\end{aligned}\right.$ (10)
and
$P(Y_{i,b}=k)=\left\{\begin{aligned} 1&-\beta,&\quad k=1\\ &\beta,&\quad k=0\end{aligned}\right.$ (11)
where $\alpha=\Phi\left(\frac{\tau-\mu}{\sigma}\right)$ and $\beta=\Phi\left(\frac{\tau+\mu}{\sigma}\right)$. By the probability functions, it is easy to deduce that
$E\left[(X_{1}-X_{2})^{2}\right]=2\sigma^{2},\qquad\quad\leavevmode\nobreak\ \leavevmode\nobreak\ E\left[(X_{1,b}-X_{2,b})^{2}\right]=2\alpha-2\alpha^{2},$
$E\left[(Y_{1}-Y_{2})^{2}\right]=2\sigma^{2},\qquad\qquad E\left[(Y_{1,b}-Y_{2,b})^{2}\right]=2\beta-2\beta^{2},$
$E\left[(X_{1}-Y_{2})^{2}\right]=2\sigma^{2}+4\mu^{2},\quad E\left[(X_{1,b}-Y_{1,b})^{2}\right]=\alpha+\beta-2\alpha\beta.$
With these equations, the discrimination $D$ of original data, as specified in Definition 2, can be further expressed as
$D=\frac{E[(X_{1}-Y_{1})^{2}]}{E[(X_{1}-X_{2})^{2}]+E[(Y_{1}-Y_{2})^{2}]}=\frac{\sigma^{2}+2\mu^{2}}{2\sigma^{2}},$ (12)
and similarly, the discrimination $D_{b}$ of binary quantized data, as specified in Definition 3, can be written as
$D_{b}=\frac{E[(X_{1,b}-Y_{1,b})^{2}]}{E[(X_{1,b}-X_{2,b})^{2}]+E[(Y_{1,b}-Y_{2,b})^{2}]}=\frac{\alpha-2\alpha\beta+\beta}{(2\alpha-2\alpha^{2})+(2\beta-2\beta^{2})}.$ (13)
Next, we are ready to prove that $D_{b}>D$ under the condition (8). By Equations (12) and (13), it is easy to see that $D_{b}>D$ is equivalent to
$(\sigma^{2}+2\mu^{2})\alpha^{2}-2(\sigma^{2}\beta+\mu^{2})\alpha+(\sigma^{2}+2\mu^{2})\beta^{2}-2\mu^{2}\beta>0.$ (14)
This inequality can be viewed as a quadratic inequality in $\alpha$, which has the discriminant:
$\Delta=4\mu^{4}+16(1-\beta)\mu^{2}\beta>0.$
By the above inequality, the inequality (14) holds when $\alpha\in(-\infty,\alpha_{1})\cup(\alpha_{2},+\infty)$, where
$\alpha_{1}=\beta+\frac{\mu^{2}(1-2\beta)-\mu\sqrt{\mu^{2}+4\beta(1-\beta)}}{1+\mu^{2}},$
and
$\alpha_{2}=\beta+\frac{\mu^{2}(1-2\beta)+\mu\sqrt{\mu^{2}+4\beta(1-\beta)}}{1+\mu^{2}}.$ (15)
Given (15), we can further derive $\alpha_{2}>\beta$, since $\mu^{2}(1-2\beta)+\mu\sqrt{\mu^{2}+4\beta(1-\beta)}>0$. However, this result contradicts the conclusion that $\alpha<\beta$ we can derive with the probability mass functions shown in (10) and (11), mainly by the increasing property of $\Phi(\cdot)$. So the solution to the inequality (14) should be $\alpha\in(-\infty,\alpha_{1})$, implying $\beta-\alpha+\frac{\mu^{2}(1-2\beta)-\mu\sqrt{\mu^{2}+4\beta(1-\beta)}}{1+\mu^{2}}>0$.
A.2 Proof of Theorem 5
Let $X_{1}$ and $X_{2}$ be i.i.d. samples of $X$, and $Y_{1}$ and $Y_{2}$ be i.i.d. samples of $Y$. Denote $X_{i,t}=f_{t}(X_{i};\tau)$ and $Y_{i,t}=f_{t}(Y_{i};\tau)$, where $i=1,2$. By the distributions of $X$ and $Y$ specified in Property 1 and the ternary quantization $f_{t}(x;\tau)$ defined in Equation (2), the probability mass functions of $X_{i,t}$ and $Y_{i,t}$ can be derived as
$P(X_{i,t}=k)=\left\{\begin{aligned} &\beta,&k&=1\\ 1-&\alpha-\beta,&k&=0\\ &\alpha,&k&=-1\end{aligned}\right.$ (16)
$P(Y_{i,t}=k)=\left\{\begin{aligned} &\alpha,&k&=1\\ 1-&\alpha-\beta,&k&=0\\ &\beta,&k&=-1\end{aligned}\right.$ (17)
where $\alpha=\Phi(\frac{-\tau-\mu}{\sigma})$ and $\beta=\Phi(\frac{-\tau+\mu}{\sigma})$.
Then, by Definition 3, the discrimination $D_{t}$ of ternary quantization can be derived as
$D_{t}=\frac{E[(X_{1,t}-Y_{1,t})^{2}]}{E[(X_{1,t}-X_{2,t})^{2}]+E[(Y_{1,t}-Y_{2,t})^{2}]}=\frac{(\alpha+\alpha^{2}-2a\beta+\beta+\beta^{2})}{2(\alpha-\alpha^{2}+2\alpha\beta+\beta-\beta^{2})}.$ (18)
By Equations (12) and (18), it can be seen that $D_{t}>D$ is equivalent to
$\frac{(\alpha+\beta)+(\alpha-\beta)^{2}}{2(\alpha+\beta)-2(\alpha-\beta)^{2}}>\frac{\sigma^{2}+2\mu^{2}}{2\sigma^{2}},$
which can simplify to
$\alpha^{2}-(2\beta+\mu^{2})\alpha+\beta^{2}-\mu^{2}\beta>0.$ (19)
Clearly, (19) can be regarded as a quadratic inequality in $\alpha$, with its discriminant:
$\Delta=\mu^{4}+8\mu^{2}\beta>0.$
This inequality implies that the inequality (19) holds when $\alpha\in(-\infty,\alpha_{1})\cup(\alpha_{2},+\infty)$, where
$\alpha_{1}=\beta+\frac{\mu^{2}-\sqrt{\mu^{4}+8\mu^{2}\beta}}{2}$
and
$\alpha_{2}=\beta+\frac{\mu^{2}+\sqrt{\mu^{4}+8\mu^{2}\beta}}{2}.$ (20)
In (20), the term $\mu^{2}+\sqrt{\mu^{4}+8\mu^{2}\beta}>0$, implying $\alpha_{2}>\beta$. In contrast, we will derive $\alpha<\beta$ by the probability functions shown in Equations (16) and (17), particularly by the increasing property of $\Phi(\cdot)$. By this contradiction, we can say that $D_{t}>D$ holds only under the case of $\alpha\in(-\infty,\alpha_{1})$, namely
$\beta-\alpha+\frac{\mu^{2}-\sqrt{\mu^{4}+8\mu^{2}\beta}}{2}>0.$
Appendix B Solution algorithms for Equations (8) and (9)
In this section, we present two algorithms for deriving the quantization thresholds $\beta$ that can enhance feature discrimination, provided they exist as specified in Equations (8) and (9) (in Theorems 4 and 5) for binary and ternary quantization, respectively.
B.1 Solution algorithm for Equation (8)
To derive a threshold $\beta$ that satisfies the inequality (8), we can minimize the objective function:
$g(\tau)=-\beta+\alpha-\frac{\mu^{2}(1-2\beta)-\mu\sqrt{\mu^{2}+4\beta(1-\beta)}}{1+\mu^{2}}$
by iteratively descending the gradient:
$g^{\prime}(\tau)=-\frac{1-\mu^{2}}{1+\mu^{2}}\beta^{\prime}+\alpha^{\prime}+\frac{\mu(2\beta^{\prime}-4\beta\beta^{\prime})}{(1+\mu^{2})\sqrt{\mu^{2}+4\beta(1-\beta)}}$
where $\alpha=\varphi\left(\frac{\tau-\mu}{\sigma}\right)$, $\beta=\varphi\left(\frac{\tau+\mu}{\sigma}\right)$, and $\varphi(\cdot)$ is the probability density function of the standard normal distribution. The gradient descent step size $\gamma$ in the $k$-th iteration can be selected with the Armijo rule (Bertsekas, 1997):
$g(\tau^{(k)}-\gamma\cdot g^{\prime}(\tau^{(k)})\leq g(\tau^{(k)})-c\cdot\gamma\cdot(g^{\prime}(\tau^{(k)}))^{2}$
where $c\in(0,1)$ is a constant.
B.2 Solution algorithm for Equation (9)
In a similar way, we can derive a threshold $\beta$ that holds the inequality (9) by minimizing
$g(\tau)=-\beta+\alpha-\frac{\mu^{2}-\sqrt{\mu^{4}+8\mu^{2}\beta}}{2}$
via iteratively descending the gradient:
$g^{\prime}(\tau)=-\beta^{\prime}+\alpha^{\prime}+\frac{2\mu\beta^{\prime}}{\sqrt{\mu^{2}+8\beta}}$
where $\alpha=\varphi\left(\frac{-\tau-\mu}{\sigma}\right)$, $\beta=\varphi\left(\frac{-\tau+\mu}{\sigma}\right)$, and $\varphi(\cdot)$ is the probability density function of the standard normal distribution. The gradient descent step size can be determined with the Armijo rule described above.
B.3 Experimental verification
Our solution algorithms are expected to produce quantization thresholds that are conductive to enhancing feature discrimination. To verify this, we evaluate the KNN classification accuracy for both binary and ternary quantization data derived from synthetic and real datasets, as described in Sections 4.1 and 4.2. In our algorithms, we set the parameter $c=10^{-3}$ for the Armijo rule and halt the iteration when $|g^{\prime}(\tau^{(k)})|_{2}<10^{-12}$ or the maximum iteration number is reached. For comparison, we also assess the performance of the conventional quantization method, which determines quantization thresholds by minimizing $\ell_{2}$-norm quantization errors (MQE) (Li et al., 2016). The experimental results are presented in Tables 1 and 2. As shown, our quantization data achieve superior classification accuracy than the conventional MQE data and the original data.
Table 1: Classification accuracy (%) of the original data, as well as the binary quantization data derived with the conventional MQE method and our method. The best results are highlighted in bold.
Dataset Original Data Binary data (MQE) Binary data (Ours)
YaleB 98.93 99.52 99.74
CIFAR10 94.08 92.58 93.54
TIMIT 92.58 94.35 94.46
Newsgroup 60.78 62.41 62.50
ImageNet 91.88 91.79 93.22
Synthetic Data 89.78 81.89 90.98
Table 2: Classification accuracy (%) of the original data, as well as the ternary quantization data derived with the conventional MQE method and our method. The best results are highlighted in bold.
Dataset Original data Ternary data (MQE) Ternary data (Ours)
YaleB 98.93 99.64 99.70
CIFAR10 94.08 96.56 96.91
TIMIT 92.58 94.10 94.15
Newsgroup 60.78 65.47 67.10
ImageNet 91.88 94.66 95.00
Synthetic Data 89.78 84.05 89.81
Appendix C Other experimental results
C.1 Numerical validation
(a) Theoretical results for the data with distribution parameters $\mu=0.99$ and $\sigma^{2}=0.02$
(b) Numerical results for the data with distribution parameters $\mu=0.99$ and $\sigma^{2}=0.02$
(c) Theoretical results for the data with distribution parameters $\mu=0.76$ and $\sigma^{2}=0.42$
(d) Numerical results for the data with distribution parameters $\mu=0.76$ and $\sigma^{2}=0.42$
Figure 8: Consider the binary quantization on two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ as specified in Property 1. For two kinds of data with distribution parameters ($\mu=0.99$, $\sigma^{2}=0.02$) and ($\mu=0.76$, $\sigma^{2}=0.42$), the values for the left and right side of Equations (8) are provided in (a) and (c) respectively; and their discrimination $D$ and $D_{b}$ statistically estimated with Equations (5) and (6) are illustrated in (b) and (d), respectively.
(a) Theoretical results for the data with distribution parameters $\mu=0.99$ and $\sigma^{2}=0.02$
(b) Numerical results for the data with distribution parameters $\mu=0.99$ and $\sigma^{2}=0.02$
(c) Theoretical results for the data with distribution parameters $\mu=0.66$ and $\sigma^{2}=0.56$
(d) Numerical results for the data with distribution parameters $\mu=0.66$ and $\sigma^{2}=0.56$
Figure 9: Consider the ternary quantization on two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ as specified in Property 1. For two kinds of data with distribution parameters ($\mu=0.99$, $\sigma^{2}=0.02$) and ($\mu=0.66$, $\sigma^{2}=0.56$), the values for the left and right side of Equations (9) are provided in (a) and (c) respectively; and their discrimination $D$ and $D_{t}$ statistically estimated with Equations (5) and (7) are illustrated in (b) and (d), respectively.
C.2 Classification on synthetic data: KNN and SVM
(a) Binary data ($\lambda=0$)
(b) Binary data ($\lambda=0.01$)
(c) Binary data ($\lambda=0.1$)
(d) Binary data ($\lambda=1$)
(e) Ternary data ($\lambda=0$)
(f) Ternary data ($\lambda=0.01$)
(g) Ternary data ($\lambda=0.1$)
(h) Ternary data ($\lambda=1$)
Figure 10: KNN (Cosine) classification accuracy for the 10,000-dimensional binary, ternary, and original data that are generated with the varying parameter $\lambda\in\{0,0.01,0.1,1\}$, which controls the data sparsity.
(a) Binary data ($\lambda=0$)
(b) Binary data ($\lambda=0.01$)
(c) Binary data ($\lambda=0.1$)
(d) Binary data ($\lambda=1$)
(e) Ternary data ($\lambda=0$)
(f) Ternary data ($\lambda=0.01$)
(g) Ternary data ($\lambda=0.1$)
(h) Ternary data ($\lambda=1$)
Figure 11: SVM classification accuracy for the 10,000-dimensional binary, ternary, and original data that are generated with the varying parameter $\lambda\in\{0,0.01,0.1,1\}$, which controls the data sparsity.
(a) 100-dim binary data
(b) 10,000-dim binary data
(c) 100-dim ternary data
(d) 10,000-dim ternary data
Figure 12: KNN (Cosine) classification accuracy for the binary, ternary, and original data generated with the sparsity parameter $\lambda=1$, and with varying dimensions $n\in\{100,10000\}$.
(a) 1-dim binary data
(b) 100-dim binary data
(c) 10,000-dim binary data
(d) 1-dim ternary data
(e) 100-dim ternary data
(f) 10,000-dim ternary data
Figure 13: SVM classification accuracy for the binary, ternary, and original data generated with the sparsity parameter $\lambda=1$, and with varying dimensions $n\in\{1,100,10000\}$.
C.3 Classification on real data: KNN and SVM
(a) Binary data with KNN (Euclidean)
(b) Binary data with SVM
(c) Ternary data with KNN (Euclidean)
(d) Ternary data with SVM
Figure 14: Classification accuracy for the binary, ternary, and original data by KNN (Euclidean distance) and SVM on CIFAR10. The parameter $\gamma$ corresponds to a threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can achieve improved or at least comparable classification performance compared to the original data.
(a) Binary quantization on YaleB
(b) Ternary quantization on YaleB
(c) Binary quantization on CIFAR10
(d) Ternary quantization on CIFAR10
(e) Binary quantization on TIMIT
(f) Ternary quantization on TIMIT
(g) Binary quantization on Newsgroup
(h) Tenary quantization on Newsgroup
Figure 15: Classification accuracy for the binary, ternary, and original data by KNN (Cosine distance) on four different datasets. The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can achieve improved classification performance compared to the original data.
(a) YaleB
(b) CIFAR10
(c) TIMIT
(d) Newsgroup
Figure 16: The histogram (blue bar) of the element values across one dimension of the feature vectors within a single class of samples selected from four different datasets, accompanied with a Gaussian fitting curve (red line).
C.4 Classification on real data: Multilayer Perceptrons (MLP) and Decision Trees
(a) Binary quantization on YaleB
(b) Ternary quantization on YaleB
(c) Binary quantization on CIFAR10
(d) Ternary quantization on CIFAR10
(e) Binary quantization on Newsgroup
(f) Ternary quantization on Newsgroup
Figure 17: MLP-based binary classification accuracy for the binary, ternary, and original data on three different datasets. The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can achieve improved or at least comparable classification performance compared to the original data.
(a) Binary quantization on YaleB
(b) Ternary quantization on YaleB
(c) Binary quantization on CIFAR10
(d) Ternary quantization on CIFAR10
(e) Binary quantization on Newsgroup
(f) Ternary quantization on Newsgroup
Figure 18: Decision trees-based binary classification accuracy for the binary, ternary, and original data on three different datasets. The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can achieve improved or at least comparable classification performance compared to the original data.
C.5 Binary and Multiclass Classifications on ImageNet1000 (with 1000 classes of samples)
(a) Binary quantization
(b) Ternary quantization
Figure 19: Binary classification accuracy for the binary, ternary, and original data in ImageNet1000, using the classifier KNN (Euclidean distance). The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can enhance the binary classification accuracy compared to the original data.
(a) Binary quantization
(b) Ternary quantization
Figure 20: Multiclass (1000-class) classification accuracy for the binary, ternary, and original data in ImageNet1000, using the classifier KNN (Euclidean distance). The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there are quantization threshold $\tau$ values that can achieve improved or at least comparable classification accuracy compared to the original data.